OpenStudy (anonymous):

Solve the equation z^4-(1+i)=0 for the complex variable z. Show Steps.

OpenStudy (perl):

z^4 = 1 + i

OpenStudy (perl):

change 1 + i to trig form or exponential form , first lets use exponential

OpenStudy (perl):

also keep in mind 1 + i corresponds to a radius is sqrt ( 1 + 1 ) = sqrt 2 , and angle is pi/4 so z^4 = sqrt 2 e^(pi/4*i)

OpenStudy (perl):

so z = ( sqrt 2 ) ^1/4 * e^ [( pi/4*i + 2pi*n) *1/4 ]

OpenStudy (perl):

we can use demoivres z^n = r [ cos (theta +2pi*k) + i sin (theta+2pi*k) ] ----> z = r^(1/n) [cos ( 1/n*theta + 2pi/k*n) + sin ( theta/n + 2pi/k*n) ]

OpenStudy (perl):

and we are given a + bi , so theta = arctan b/a , and r = sqrt ( a^2 + b^2)

OpenStudy (perl):